# Deliberative XAI: How Explanations Impact Understanding and Decision-Making of AI Novices in Collective and Individual Settings

### v1

@@ 5.2. Benefits and challenges of using explanations in collaborative settings
Challenges: Explanations must adapt to different groups and require constructive dynamics. While our explanations were effective in supporting interaction and opinion formation in most focus groups, we also observed several challenges resulting from group compositions and trust between participants. While the explanations typically worked well in groups of domain experts, especially G (executive staff) and H (job trainers and counselors), special interaction situations occurred in Groups B and H due to two participants’ leading roles in the organizations. This could have resulted in disrupted interaction due to fear of being evaluated by someone perceived as superordinate, but, in fact, had no negative impact on collaboration. In Groups E and F (decision subjects), however, the unfamiliarity of participants and the explanation’s high access threshold led to a lack of interactions and discussion, indicating that the explanation did not fully support these compositions. In these groups, a lack of trust between participants likely disrupted interactions when participants were unknown to each other, amplifying effects such as the fear of being evaluated (Nokes-Malach et al., 2015). This underscores the importance of creating trust between group members in collaborative XAI settings (Johnson and Johnson, 1985). Intuitive measures could be the introduction of a simple task that the group solves collaboratively before engaging with explanations, such as the Wason card selection task (Wason, 1968), or creating an atmosphere that allows participants to interact in roles (e.g., proponents and opposition) to facilitate discussion without revealing personal opinions. Future work should examine how such measures can be incorporated into explanation design to support interaction in groups of comparable compositions. Lastly, and surprisingly, we did not observe many cognitive biases such as ”groupthink” (Janis, 1972) (aiming for consensus at the cost of negotiation and alternatives) or the ”equality bias” (Naiseh et al., 2024) (downplaying one’s expertise to weigh everyone’s opinion) in our study. In one example of groupthink, participants in Group A changed their vote to ”Yes” to reach a group decision, perhaps driven by their (stated) lack of being personally affected by the system’s deployment. However, this is contrasted by participants in Group D, who debated at length about the system’s deployment without reaching a consensus, despite also not being affected personally. Potential measures to avoid groupthink in discussion could thus be to encourage debate, which again could be the introduction of roles to improve the ”dialectic argumentation” (Mercier and Sperber, 2011), and to explain the system in a way that makes it more personally relevant to participants (Wiggins and McTighe, 2005), e.g., by emphasizing connections to their own experiences.
5.3. Explanation design suggestions
@@
We summarize the implications of our findings in the form of suggestions for the design of explanations that are suited to AI novices in individual and group settings, as depicted in Figure 7.
@@ 6. Limitations
In this section, we briefly describe limitations of our study that the reader should consider. Due to the limited sample size, we did not analyze the impact of sex and/or gender on our results, limiting the results’ generalizability regarding these aspects but not their overall validity. Further, our participants were recruited from organizations and networks in the same geographical region, perhaps resulting in regional or cultural biases. The presented use case is embedded in a specific sociotechnical context (Ehsan et al., 2023a) that might affect participants’ understanding and perceptions (e.g., perceptions might differ between employability prediction and credit approval), and thus a change in the domain might also change the explanations’ effect. However, this does not limit the transferability of the explanation design, which can be seen as a template that can be adjusted to other use cases. We further note that our participant sample has a bias toward university education in the single interviews, which we addressed by comparing these participants mostly with university-educated participants in the focus groups.
@@ 7. Conclusion
In this paper, we proposed and empirically validated a question-driven, modular explanation design for individual and collaborative settings. We conducted an interview study with AI novices in 8 focus groups and 12 single interviews, analyzing the explanations’ effect on understanding and decision confidence, participants’ perceptions of key information, and the interaction processes in both settings. We found that explanations were effective at supporting and calibrating participants’ understanding and at informing individual and collective decision-making. Our findings further showed that our explanations support focused and detail-oriented interaction in individual settings while also being suited to be shared and discussed in group settings. In groups with constructive dynamics, interactions with the explanations led to ”collaborative success” (Nokes-Malach et al., 2015) and informed group discussions about pivotal topics regarding the system’s deployment. For groups in which the explanations were less effective, we suggest adjusting the group dynamic, modifying the explanation’s format, and using co-design approaches to better adapt explanations to various needs. Based on these findings, we present suggestions for the design of explanations for AI novices and outline how their application can support collaborative reasoning and discussion. With this work, we aim to contribute an explanation design approach to the field of collaborative XAI and present implications for how explanations can support stakeholder deliberation about public AI systems.
@@ Acknowledgements.
- Burns et al. (2023) Alyxander Burns, Christiana Lee, Ria Chawla, Evan Peck, and Narges Mahyar. 2023. Who Do We Mean When We Talk About Visualization Novices?. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (Hamburg, Germany) (CHI ’23). Association for Computing Machinery, New York, NY, USA, Article 819, 16 pages. https://doi.org/10.1145/3544548.3581524
@@
- European Commission (2021) European Commission. 2021. Laying Down Harmonised Rules on Artificial Intelligence. digital-strategy.ec.europa.eu/en/library/proposal-regulation-laying-down-harmonised-rules-artificial-intelligence.
@@
- Lee et al. (2019) Min Kyung Lee, Daniel Kusbit, Anson Kahng, Ji Tae Kim, Xinran Yuan, Allissa Chan, Daniel See, Ritesh Noothigattu, Siheon Lee, Alexandros Psomas, and Ariel D. Procaccia. 2019. WeBuildAI: Participatory Framework for Algorithmic Governance. Proceedings of the ACM on Human-Computer Interaction 3, CSCW (Nov. 2019), 1–35. https://doi.org/10.1145/3359283
@@
- Shen et al. (2020) Hong Shen, Haojian Jin, Angel Alexander Cabrera, Adam Perer, Haiyi Zhu, and Jason I. Hong. 2020. Designing Alternative Representations of Confusion Matrices to Support Non-Expert Public Understanding of Algorithm Performance. Proceedings of the ACM on Human-Computer Interaction 4, CSCW2 (Oct. 2020), 1–22. https://doi.org/10.1145/3415224
@@
- Yurrita et al. (2022) Mireia Yurrita, Dave Murray-Rust, Agathe Balayn, and Alessandro Bozzon. 2022. Towards a multi-stakeholder value-based assessment framework for algorithmic systems. In Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency (Seoul, Republic of Korea) (FAccT ’22). Association for Computing Machinery, New York, NY, USA, 535–563. https://doi.org/10.1145/3531146.3533118
@@
- Zhang et al. (2023) Angie Zhang, Olympia Walker, Kaci Nguyen, Jiajun Dai, Anqing Chen, and Min Kyung Lee. 2023. Deliberating with AI: Improving Decision-Making for the Future through Participatory AI Design and Stakeholder Deliberation. Proceedings of the ACM on Human-Computer Interaction 7, CSCW1 (April 2023), 1–32. https://doi.org/10.1145/3579601
@@
Appendix
The supplementary material provides additional information on the study and use case. In Section A, we provide the features used by the employment prediction algorithm, the mock newspaper article introducing the use case to participants, and the study task description. Section B lists the self-report and interview questions that were asked during the first and second self-reports.

### v2

@@ 5.1. Do AI novices learn and deliberate about AI better together or individually?
The importance of the social dynamic became evident in groups where members were not familiar and had trouble engaging with the explanations. In groups G and H, for example, the social dynamic bridged understanding issues of individual participants and acted against discouragement. In groups E and F, in contrast, these understanding issues eventually led participants to abandon understanding, as the social atmosphere did not support them in overcoming them. Here, a lack of trust or simply unfamiliarity between participants likely amplified effects such as social loafing and the fear of being evaluated (Nokes-Malach et al., 2015). This underscores the importance of creating trust between group members in collaborative XAI settings (Johnson and Johnson, 1985). Intuitive measures could be the introduction of a simple task that the group solves collaboratively before engaging with explanations, such as the Wason card selection task (Wason, 1968). Another measure could be the introduction of roles (e.g., proponents and opposition), as has been done with the “devil’s advocate” in previous work (Chiang et al., 2024), to facilitate discussion and close examination of the proposal. Future work should examine how such measures can be incorporated into explanation design to support interaction in groups of comparable compositions. Lastly, regarding cognitive biases, we observed an effect resembling some aspects of “groupthink” (Janis, 1971) when participants in Group A changed their vote to “Yes” to reach a group decision. We argue that this effect originated in the lack of detailed interaction with explanations and, possibly, a perceived low degree of personal affection by the system’s deployment. However, this is contrasted by participants in Group D, who debated at length about the system’s deployment without reaching a consensus, despite not being directly affected. Potential measures to avoid groupthink in discussion could thus be to encourage debate, which again could be the introduction of roles to improve the “dialectic argumentation” (Mercier and Sperber, 2011), and to explain the system in a way that makes it more personally relevant to participants (Wiggins and McTighe, 2005), e.g., by emphasizing connections to their own experiences.
5.2. Did the explanations improve participants’ understanding?
In Section 4.1.5, we described that the explanations helped participants develop different ‘facets of understanding’ (Wiggins and McTighe, 2005). In groups, participants were encouraged to ‘explain’ information to each other and ‘empathize’ with others’ experiences, while individuals could better ‘apply’ information in the study tasks. We further described that groups’ interactions with explanations realized mechanisms of “collaborative success” (Nokes-Malach et al., 2015). We thus conclude that the explanations had a positive effect on understanding. However, a more complete answer requires that we consider the difference between measurement methods and true cognitive states. In Section 4.1, we described that a majority of participants reported unchanged understanding after the explanation phase (Table 3) but, paradoxically, described verbally that their understanding improved; two seemingly incongruent pieces of evidence. We explain these contradictory findings with a process we call calibrating understanding. The term describes that participants tend to report understanding not in absolute terms, or even in relation to past understanding, but in relation to the currently available information. Participants explicitly stated that they calibrated their interpretation of ’good understanding’ according to their knowledge of the information basis, which differed before and after the explanation phase. The calibration process can be traced by using concepts from the cognitive sciences: Participants i) reported their initial understanding after reading the use case introduction, they then ii) saw the explanations and realized that they had understanding gaps (Rozenblit and Keil, 2002), which they iii) proceeded to locate and close (Keil, 2006), however, they iv) also realized that they could not look at every available explanation and would develop at most a “partial understanding” (Keil, 2019), which they v) rated accordingly in the second self-reports.
Previous studies in XAI documented similar effects caused by white-box explanations (Cheng et al., 2019), which, due to their high information density, led to increased ‘objective’ understanding but decreased self-reported understanding. Importantly, (Papenmeier et al., 2022) found similar discrepancies when participants who received no explanation gave higher understanding scores than participants who received faithful explanations; a discrepancy our findings might explain. In the same study, similar discrepancies also occurred between trust self-reports and observations of behavior (Papenmeier et al., 2022). In line with these findings, we argue that the calibration effect should be accounted for when measuring understanding, for example, by eliciting an additional metric that captures the perceived scope of available information. A potential reporting question could be “How much information do you feel you currently have about the presented AI system?”, combined with a 5-point scale ranging from “very little” to “very much”. Self-reported understanding could then be compared with self-reported information scope and verbal responses to acquire a more complete picture. Recent work in XAI has further proposed understanding measurement based on participants’ abilities (Speith et al., 2024). This approach appears promising, as the ability to calculate study task 2 was a relevant metric in our study. We thus see eliciting understanding via multiple measures and exploring how these measures can be combined in individual and group settings as a direction for future research.
5.3. Which real-world settings would benefit from explainable AI in groups?
In Section 2.3.2, we described several settings where citizens gather to discuss and form opinions on matters of public interest. These included referendums, forums, and community-based spaces. This paper investigates settings suitable for deliberating the deployment of public AI systems, an issue that we frame as a matter of public interest due to the scope and severity of its potential consequences. Having established that using explanations in group settings benefits participants’ understanding, decision confidence, and decision-making processes, it is worthwhile to consider how this setting could be employed in real-world contexts. One answer can be given based on participants’ feedback, who stated that training in their job agency should employ a similar format to educate about AI. Notably, this feedback was given by domain experts (Groups C, G) and decision subjects (Group E), suggesting that the setting would suit both stakeholder groups for an educational intervention. Similarly, P3 explained that the explanation approach could be helpful if a similar system were used in their care facility by embedding it in the team’s regular meetings, in which difficult cases are discussed and joint decisions are made. These insights are in line with previous work on XAI in public institutions. Notably, Lee et al. (2019b) and Weitz et al. (2024) conducted participatory workshops to design explanations with end-users in the public sector, finding that co-designing explainable AI helps in considering the needs of both clients and end-users. We envision that collaborative settings and ‘mini-publics’ (Fung, 2003) could be useful in many contexts that aim to strengthen participatory democracy with respect to AI. Potential areas of application could be professional consultation workshops for citizens affected by algorithmic decisions, comparable to legal clinics (Ontario, 2025), community-based education and training interventions, such as “contestation cafés” (Collins et al., 2024), or union forums that inform and organize employees’ voices about the use of AI in their institution (Kaur et al., 2022). On a different note, Crivellaro et al. (2019) found that participatory formats that aim to connect communities to public institutions can suffer from a lack of crucial information (e.g., budgets), which could be alleviated by an information structure such as the presented explanation design. In short, we propose that explainable AI in collective settings could be a valuable engagement format for contexts in which public AI could impact people’s lives. Future work could explore how collective XAI settings could be implemented in these contexts as part of responsible AI initiatives and in connection to both institutionalized (Costanza-Chock et al., 2022) and user-based (Shen et al., 2021) auditing practices.
5.4. What’s missing from the explanation design and how could it be improved?
In Sections 4.1.4, we described that participants appreciated the explanations’ comprehensive and flexible information selection and self-directed and active exploration. However, they also noted that the explanations have a high access threshold and require adjustment to the modular structure, making oversight difficult. A digital version of the explanation design could improve the overview through summaries and navigation while allowing for simple language options and cross-references. As the simple awareness of the scope of information also seemed to overwhelm participants, approaches to condense the scope and selection would be beneficial. An example could be a recommendation system that would suggest to participants explanations from different categories and levels of detail based on their stated interests and technical knowledge. However, we also emphasize that identifying explanation subtopics and splitting them up into levels of detail is challenging. Our explanations’ structure provides different levels of soundness and completeness (Kulesza et al., 2013), but deciding how the available information is allocated into this structure requires a subjective choice. In our study, detail levels 2 and 3 were supposed to convey more detailed but also more difficult explanations than the base levels. Still, some participants stated that the most critical information for them resided in level 3. We see avenues for future work in selecting and hierarchically structuring information to be included in explanations for AI systems, such that they allow for exploration while not obscuring essential information. This essential information should balance textual and visual design to ensure its uptake does not rely only on textual understanding.
Due to the modular structure and the complexity of some of the information, several participants used intuition in answering the study tasks instead of truly searching for the explanations for solutions. Framing this in the dual-process theory of cognition, we observed that participants in these moments used System 1 (intuitive heuristics) rather than System 2 thinking (analytical reasoning) (Buçinca et al., 2021; Kahneman, 2011). Lambe et al. (2016) listed strategies to counteract this tendency and encourage analytical thinking in the medical domain, including checklists, cognitive forcing mechanisms (consideration of alternative diagnoses, reconsideration of diagnoses), guided reflection, and use of particular reasoning approaches. Buçinca et al. (2021) further tested cognitive forcing mechanisms in an AI-assisted decision-making scenario, in which they used three interventions: showing participants an AI decision only on demand, showing the AI decision only after the participants made their own predictions, and letting them wait before showing them the AI decision. These mechanisms improved the performance of the human-AI teams but led users to dislike the interface’s usability, presenting a trade-off. Nonetheless, we concur that these strategies should be considered in future explanation design to “ensure that people will exert effort to attend to those explanations” (Buçinca et al., 2021).
5.5. Summary of explanation design suggestions
@@
We summarize the implications of our findings in the form of suggestions for the design of explanations suited to AI novices in individual and group settings in Figure 8.
@@ 6. Limitations
Like any research, this study had limitations. Due to the limited sample size, we did not analyze the impact of sex and/or gender on our results, limiting the results’ generalizability regarding these aspects but not their overall validity. Further, our participants were recruited from organizations and networks in the same geographical region, perhaps resulting in regional or cultural biases. The presented use case is further embedded in a specific sociotechnical context (Ehsan et al., 2023a) that might affect participants’ understanding and perceptions (e.g., perceptions might differ between employability prediction and credit approval), and thus, a change in the domain might also change the explanations’ effect. However, this does not limit the transferability of the explanation design, which can be seen as a template that can be adjusted to other use cases. We further note that our participant sample is biased toward university education in the single interviews, which we addressed by comparing these participants mostly with university-educated participants in the focus groups.
We are further aware that the cooperation with civil society organizations in the recruitment of participants could have led to selection biases, especially in the form of convenience sampling (over-representation of readily available participants), self-selection bias (over-representation of strongly motivated participants), and interviewer bias (over-representation of agreeable or compatible participants) (Collier, 1995). We aimed to counteract these biases by defining research goals and methods clearly before recruiting participants, by using multiple recruitment sources and methods, ensuring that group composition was diversified, and by reflecting on possible sampling influences in the analysis of results. Due to the qualitative approach, our recruitment strategy further did not aim for statistical generalizability but instead intended to cover a variety of “theoretically relevant cases” (Collier, 1995) and “careful contextualization” (Collier and Mahoney, 1996) to examine our research questions.
@@ 7. Conclusion
This paper tested a question-driven, modular explanation design with AI novices in groups and individual settings. We conducted an interview study involving 8 focus groups and 12 single interviews. We analyzed them to examine the effect of explanations on understanding, decisions, and decision confidence, participants’ perceptions of key information, and the interaction processes in both settings. We found that explanations supported participants’ understanding and decision-making differently, encouraging focused interaction in individual settings and shared understanding in group settings. Even though individuals could not exchange with others, the explanations still led to increased decision confidence and changes by supporting internal deliberation. In groups, the explanation design afforded a set of interactions that allowed participants to support each other’s understanding, and further provided grounds for exchanging arguments about key aspects regarding the system’s deployment. For groups that experienced collaborative failure, we suggest the modification of the explanation’s design to highlight essential information and measures to create a more productive social dynamic. With this work, we aim to showcase the potential of combining explanations with group settings to enable AI novices to understand and deliberate about public AI systems.
@@ Acknowledgements.
- Baron (2005) Robert S. Baron. 2005. So Right It’s Wrong: Groupthink and the Ubiquitous Nature of Polarized Group Decision Making. In Advances in Experimental Social Psychology. Vol. 37. Elsevier, 219–253. https://doi.org/10.1016/S0065-2601(05)37004-3
@@
- Biran and McKeown (2017) Or Biran and Kathleen McKeown. 2017. Human-Centric Justification of Machine Learning Predictions. In Proceedings of the Twenty-Sixth International Joint Conference on Artificial Intelligence, IJCAI-17. 1461–1467. https://doi.org/10.24963/ijcai.2017/202
@@
- Buçinca et al. (2021) Zana Buçinca, Maja Barbara Malaya, and Krzysztof Z. Gajos. 2021. To Trust or to Think: Cognitive Forcing Functions Can Reduce Overreliance on AI in AI-assisted Decision-making. Proceedings of the ACM on Human-Computer Interaction 5, CSCW1 (April 2021), 1–21. https://doi.org/10.1145/3449287
@@
- Carter et al. (2014) Nancy Carter, Denise Bryant-Lukosius, Alba DiCenso, Jennifer Blythe, and Alan J. Neville. 2014. The Use of Triangulation in Qualitative Research. Oncology Nursing Forum 41, 5 (Sept. 2014), 545–547. https://doi.org/10.1188/14.ONF.545-547
@@
- Chiang et al. (2023) Chun-Wei Chiang, Zhuoran Lu, Zhuoyan Li, and Ming Yin. 2023. Are Two Heads Better Than One in AI-Assisted Decision Making? Comparing the Behavior and Performance of Groups and Individuals in Human-AI Collaborative Recidivism Risk Assessment. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems. ACM, Hamburg Germany, 1–18. https://doi.org/10.1145/3544548.3581015
- Chiang et al. (2024) Chun-Wei Chiang, Zhuoran Lu, Zhuoyan Li, and Ming Yin. 2024. Enhancing AI-Assisted Group Decision Making through LLM-Powered Devil’s Advocate. In Proceedings of the 29th International Conference on Intelligent User Interfaces. ACM, Greenville SC USA, 103–119. https://doi.org/10.1145/3640543.3645199
@@
- Collier (1995) David Collier. 1995. Translating quantitative methods for qualitative researchers: The case of selection bias. American Political Science Review 89, 2 (1995), 461–466.
- Collier and Mahoney (1996) David Collier and James Mahoney. 1996. Insights and Pitfalls: Selection Bias in Qualitative Research. World Politics 49, 1 (Oct. 1996), 56–91. https://doi.org/10.1353/wp.1996.0023
- Collins et al. (2024) Robert Patrick Collins, Johan Redström, and Marco Rozendaal. 2024. The right to contestation: Towards repairing our interactions with algorithmic decision systems. (2024). https://doi.org/10.57698/V18I1.06 Publisher: International Journal of Design.
@@
- Costanza-Chock et al. (2022) Sasha Costanza-Chock, Inioluwa Deborah Raji, and Joy Buolamwini. 2022. Who Audits the Auditors? Recommendations from a field scan of the algorithmic auditing ecosystem. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 1571–1583. https://doi.org/10.1145/3531146.3533213
- Crivellaro et al. (2019) Clara Crivellaro, Rob Anderson, Daniel Lambton-Howard, Tom Nappey, Patrick Olivier, Vasilis Vlachokyriakos, Alexander Wilson, and Pete Wright. 2019. Infrastructuring Public Service Transformation: Creating Collaborative Spaces between Communities and Institutions through HCI Research. ACM Trans. Comput.-Hum. Interact. 26, 3, Article 15 (May 2019), 29 pages. https://doi.org/10.1145/3310284
@@
- Eslami et al. (2016) Motahhare Eslami, Karrie Karahalios, Christian Sandvig, Kristen Vaccaro, Aimee Rickman, Kevin Hamilton, and Alex Kirlik. 2016. First I ”like” It, Then I Hide It: Folk Theories of Social Feeds. In Proceedings of the 2016 CHI Conference on Human Factors in Computing Systems (San Jose, California, USA) (CHI ’16). Association for Computing Machinery, New York, NY, USA, 2371–2382. https://doi.org/10.1145/2858036.2858494
- European Commission (2024) European Commission. 2024. Laying Down Harmonised Rules on Artificial Intelligence and Amending Regulations. .
@@
- Fung (2003) Archon Fung. 2003. Survey Article: Recipes for Public Spheres: Eight Institutional Design Choices and Their Consequences. Journal of Political Philosophy 11, 3 (2003), 338–367. https://doi.org/10.1111/1467-9760.00181 arXiv:https://onlinelibrary.wiley.com/doi/pdf/10.1111/1467-9760.00181
- Gamper et al. (2020) Jutta Gamper, Günter Kernbeiß, and Michael Wagner-Pinter. 2020. Das Assistenzsystem AMAS: Zweck, Grundlagen, Anwendung. https://www.ams-forschungsnetzwerk.at/downloadpub/2020_Assistenzsystem_AMAS-dokumentation.pdf
@@
- Habermas (1991) Jurgen Habermas. 1991. The structural transformation of the public sphere: An inquiry into a category of bourgeois society. MIT press.
- Hajer and Wagenaar (2003) Maarten A. Hajer and H. Wagenaar (Eds.). 2003. Deliberative policy analysis: understanding governance in the network society. Cambridge University Press, Cambridge, UK ; New York, USA.
- Henin and Le Métayer (2022) Clément Henin and Daniel Le Métayer. 2022. Beyond explainability: justifiability and contestability of algorithmic decision systems. AI & SOCIETY 37, 4 (Dec. 2022), 1397–1410. https://doi.org/10.1007/s00146-021-01251-8
@@
- Holl et al. (2018) Jürgen Holl, Günter Kernbeiß, and Michael Wagner-Pinter. 2018. Das AMS-Arbeitsmarktchancen-Modell.
- Innes and Booher (2003) Judith E. Innes and David E. Booher. 2003. Collaborative policymaking: governance through dialogue. In Deliberative Policy Analysis (1 ed.), Maarten A. Hajer and Hendrik Wagenaar (Eds.). Cambridge University Press, 33–59. https://doi.org/10.1017/CBO9780511490934.003
- Janis (1971) IL Janis. 1971. Groupthink/Janis. IL//Psychology Today 5 (1971), 6.
@@
- Kahneman (2011) Daniel Kahneman. 2011. Thinking, fast and slow. macmillan.
- Karadzhov et al. (2023) Georgi Karadzhov, Tom Stafford, and Andreas Vlachos. 2023. DeliData: A Dataset for Deliberation in Multi-party Problem Solving. Proceedings of the ACM on Human-Computer Interaction 7, CSCW2 (Sept. 2023), 1–25. https://doi.org/10.1145/3610056
- Kaur et al. (2022) Harmanpreet Kaur, Eytan Adar, Eric Gilbert, and Cliff Lampe. 2022. Sensible AI: Re-imagining Interpretability and Explainability using Sensemaking Theory. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 702–714. https://doi.org/10.1145/3531146.3533135
@@
- Kerr and Tindale (2004) Norbert L. Kerr and R. Scott Tindale. 2004. Group Performance and Decision Making. Annual Review of Psychology 55, 1 (Feb. 2004), 623–655. https://doi.org/10.1146/annurev.psych.55.090902.142009
@@
- Lambe et al. (2016) Kathryn Ann Lambe, Gary O’Reilly, Brendan D Kelly, and Sarah Curristan. 2016. Dual-process cognitive interventions to enhance diagnostic reasoning: a systematic review. BMJ quality & safety 25, 10 (2016), 808–820.
@@
- Lee et al. (2019a) Min Kyung Lee, Anuraag Jain, Hea Jin Cha, Shashank Ojha, and Daniel Kusbit. 2019a. Procedural justice in algorithmic fairness: Leveraging transparency and outcome control for fair algorithmic mediation. Proceedings of the ACM on Human-Computer Interaction 3, CSCW (2019), 1–26.
- Lee et al. (2019b) Min Kyung Lee, Daniel Kusbit, Anson Kahng, Ji Tae Kim, Xinran Yuan, Allissa Chan, Daniel See, Ritesh Noothigattu, Siheon Lee, Alexandros Psomas, and Ariel D. Procaccia. 2019b. WeBuildAI: Participatory Framework for Algorithmic Governance. Proceedings of the ACM on Human-Computer Interaction 3, CSCW (Nov. 2019), 1–35. https://doi.org/10.1145/3359283
@@
- Liao et al. (2021) Q. Vera Liao, Milena Pribić, Jaesik Han, Sarah Miller, and Daby Sow. 2021. Question-Driven Design Process for Explainable AI User Experiences. arXiv:https://arxiv.org/abs/2104.03483 [cs.HC]
@@
- Lupia (2023) Arthur Lupia. 2023. By Design: How People Adapt to Cognitive Limitations in Politics. Topics in Cognitive Science (Sept. 2023), tops.12690. https://doi.org/10.1111/tops.12690
- Mair et al. (2019) David Mair, Laura Smillie, Giovanni La Placa, Florian Schwendinger, Milena Raykovska, Zsuzsanna Pasztor, René van Bavel, and European Commission (Eds.). 2019. Understanding our political nature: how to put knowledge and reason at the heart of political decision-making. Publications Office, Luxembourg. https://doi.org/10.2760/374191
- Mandviwalla and Olfman (1994) Munir Mandviwalla and Lorne Olfman. 1994. What do groups need? A proposed set of generic groupware requirements. ACM Trans. Comput.-Hum. Interact. 1, 3 (Sept. 1994), 245–268. https://doi.org/10.1145/196699.196715
@@
- Mercier and Landemore (2012) Hugo Mercier and Hélène Landemore. 2012. Reasoning Is for Arguing: Understanding the Successes and Failures of Deliberation. Political Psychology 33, 2 (April 2012), 243–258. https://doi.org/10.1111/j.1467-9221.2012.00873.x
@@
- Mohseni et al. (2020) Sina Mohseni, Niloofar Zarei, and Eric D. Ragan. 2020. A Multidisciplinary Survey and Framework for Design and Evaluation of Explainable AI Systems. arXiv:1811.11839 [cs.HC] https://arxiv.org/abs/1811.11839
@@
- Ontario (2025) Legal Aid Ontario. 2025. What is a legal clinic? - Legal Aid Ontario — legalaid.on.ca. https://www.legalaid.on.ca/faq/what-is-a-legal-clinic/. [Accessed 03-04-2025].
- Panigutti et al. (2023) Cecilia Panigutti, Ronan Hamon, Isabelle Hupont, David Fernandez Llorca, Delia Fano Yela, Henrik Junklewitz, Salvatore Scalzo, Gabriele Mazzini, Ignacio Sanchez, Josep Soler Garrido, and Emilia Gomez. 2023. The Role of Explainable AI in the Context of the AI Act. In Proceedings of the 2023 ACM Conference on Fairness, Accountability, and Transparency (Chicago, IL, USA) (FAccT ’23). Association for Computing Machinery, New York, NY, USA, 1139–1150. https://doi.org/10.1145/3593013.3594069
- Papenmeier et al. (2022) Andrea Papenmeier, Dagmar Kern, Gwenn Englebienne, and Christin Seifert. 2022. It’s Complicated: The Relationship between User Trust, Model Accuracy and Explanations in AI. ACM Trans. Comput.-Hum. Interact. 29, 4, Article 35 (March 2022), 33 pages. https://doi.org/10.1145/3495013
@@
- Shen et al. (2021) Hong Shen, Alicia DeVos, Motahhare Eslami, and Kenneth Holstein. 2021. Everyday Algorithm Auditing: Understanding the Power of Everyday Users in Surfacing Harmful Algorithmic Behaviors. Proceedings of the ACM on Human-Computer Interaction 5, CSCW2 (Oct. 2021), 1–29. https://doi.org/10.1145/3479577
@@
- Speith (2022) Timo Speith. 2022. A Review of Taxonomies of Explainable Artificial Intelligence (XAI) Methods. In 2022 ACM Conference on Fairness, Accountability, and Transparency. ACM, Seoul Republic of Korea, 2239–2250. https://doi.org/10.1145/3531146.3534639
@@
- Stromer-Galley (2007) Jennifer Stromer-Galley. 2007. Measuring Deliberation’s Content: A Coding Scheme. Journal of Deliberative Democracy 3, 1 (July 2007). https://doi.org/10.16997/jdd.50
@@
- Swiss Confederation (2025) Swiss Confederation. 2025. The referendum. ch.ch/en/votes-and-elections/referendum.
- Szigetvari (2018) András Szigetvari. 2018. AMS bewertet Arbeitslose künftig per Algorithmus. Der Standard (2018). https://www.derstandard.at/story/2000089095393/ams-bewertet-arbeitslose-kuenftig-per-algorithmus
@@
- Veenman et al. (2006) Marcel V. J. Veenman, Bernadette H. A. M. Van Hout-Wolters, and Peter Afflerbach. 2006. Metacognition and learning: conceptual and methodological considerations. Metacognition and Learning 1, 1 (April 2006), 3–14. https://doi.org/10.1007/s11409-006-6893-0
@@
Supplementary material
Better Together? The Role of Explanations in Supporting Novices in Individual
and Collective Deliberations about AI
The supplementary material provides additional information on the study and use case. In Section A, we provide the features used by the employment prediction algorithm, the mock newspaper article introducing the use case to participants, and the study task description. Section B lists the self-report and interview questions that were asked during the first and second self-reports. Section C depicts all explanations that participants could receive in the study. Lastly, Section D provides the complete code book, listing themes and codes generated from the qualitative analysis.